### v2

# Nothing Conformal about Adaptive Conformal Inference

Abstract
Conformal prediction is a widely-used framework for distribution-free uncertainty quantification, which generates valid prediction sets at a user-defined significance level. However, this framework relies on the assumption that the data-generating distribution is exchangeable, a condition that is frequently violated in time-series and other structured data. In such cases, the validity guarantees of conformal prediction break down.
Adaptive conformal inference (ACI) has been proposed as a solution for non-exchangeable data by dynamically adjusting the significance level to retain at least finite sample guarantees on the marginal coverage error rate. This paper demonstrates that, despite its name, ACI does not strictly require the use of conformal predictors. Instead, it can operate effectively with the more general concept of a confidence predictor, which is often computationally simpler. The key requirement is that larger significance levels correspond to smaller prediction sets, a property known as nested prediction sets.
Through experiments on synthetic and real-world data, we investigate whether ACI with conformal predictors offers advantages over confidence predictors. Our results indicate that confidence predictors can perform just as well, and sometimes better than conformal predictors in some cases, although further empirical studies are needed to determine when one approach may be preferable.
keywords:
Prediction Intervals , Coverage Guarantee , Adaptive Conformal Inference , Conformal Prediction
journal: Pattern Recognition
[…]
[inst1]organization=Department of Mathematics, Stockholm University,city=Stockholm, country=Sweden
[…]
[inst2]organization=Department of Computing, Jönköping School of Engineering,city=Jönköping, country=Sweden
1 Introduction
Conformal prediction (CP) [1, 2] is a general framework for distribution-free uncertainty quantification. It can be used with essentially any machine learning algorithm and has guaranteed properties of validity, provided that the data is drawn from a probability distribution that is exchangeable, meaning essentially that any permutation is equally probable. For details on exchangeability, see e.g. [2]. The strongest notions of validity are achieved in the online setting, where Reality outputs successive examples $z_{n}:=(x_{n},y_{n})\in\boldsymbol{Z}:=\boldsymbol{X}\times\boldsymbol{Y}$, each one consisting of an object $x_{n}\in\boldsymbol{X}$ and its associated label $y_{n}\in\boldsymbol{Y}$, CP produces prediction sets, $\Gamma_{n}^{\varepsilon}$, at a user specified significance level $\varepsilon$, using $z_{1},\dots,z_{n-1}$ and $x_{n}$ as input.
While powerful and general, the validity guarantees of CP relies on the exchangeability assumption, which is a limitation e.g. in time-series forecasting. To address these limitations, many authors have suggested solutions to retain at least some validity guarantees even when exchangeability is violated. These include introducing weights to attempt to deal with distribution shifts [3, 4], and what could be described as control theoretic approaches that attempt to achieve the desired error rate by means of varying the significance level [5, 6].
Adaptive conformal inference (ACI) [5] falls into the second category, of control theoretic approaches, and has become a popular choice for non-exchangeable data. It has been suggested as a good method for time-series forecasting [7], and the software package MAPIE [8] has implemented ACI for this purpose. The more complicated conformal PID algorithm [6] can be viewed as a generalisation of ACI, as the authors recover it as a special case of conformal PID.
In this paper, we will show that there is no need to use ACI with a conformal predictor. The same guaranteed error rates hold even if we use the more general notion of a confidence predictor, which can often be much more computationally tractable. Since nothing is gained by using the heavier machinery of conformal predictors in terms of error rate, the question is whether CP offers more efficient prediction sets.
The rest of this paper is organised as follows. Section 2 introduces conformal and confidence predictors, as well as the relevant notation. Readers familiar with the terminology used in [2] may find it convenient to skip ahead to Section 3, where ACI and its finite sample coverage guarantee is introduced, as well as a trivial but technically necessary extension of conformal and confidence predictors. Our main result is given in Section 4, where we restate Lemma 4.1 in [5], which is the key result that is used to prove the finite sample guarantee of ACI. We indicate that this lemma does not rely on any particular property of conformal predictors, or even confidence predictors. We then argue in Section 5 that, while not strictly necessary for the finite sample guarantee of ACI, confidence predictors represent the natural way of predicting at a confidence level. Our numerical experiments on synthetic and real data are described in Section 6, and Section 7 concludes. Additional details on some algorithms used in the experiments, and a summary of the results without using ACI (i.e. at a fixed significance level) are given in the appendices.
2 Theoretical background
This section introduces the relevant theoretical background on mathematical objects and notation. Readers familiar with conformal prediction, and the notation used in [2] may find it convenient to skip ahead to the next section, where ACI is introduced.
2.1 Confidence prediction in the online setting
Given two measurable spaces $\boldsymbol{X}$ and $\boldsymbol{Y}$, the former called the object space and the latter is the label space, we assume that Reality outputs successive pairs
$(x_{1},y_{1}),(x_{2},y_{2}),\dots$ (1)
called examples. For notational convenience, we write $z_{i}:=(x_{i},y_{i})\in\boldsymbol{X}\times\boldsymbol{Y}:=\boldsymbol{Z}$. The measurable space $\boldsymbol{Z}$ is called the example space. Thus, the infinite sequence (1) is an element of the measurable space $\boldsymbol{Z}^{\infty}$, and we assume that it is drawn from some probability distribution $P$ on $\boldsymbol{Z}^{\infty}$. The standard assumption in CP is that $P$ is exchangeable, but in this paper we will make no such assumption, and $P$ can be any distribution.
Definition 1 (Confidence predictor).
A confidence predictor is a measurable function
$\Gamma:\boldsymbol{Z}^{*}\times\boldsymbol{X}\times(0,1)\to 2^{\boldsymbol{Y}}$ (2)
that, for each significance level $\varepsilon\in(0,1)$ outputs the prediction set
$\Gamma_{n}^{\varepsilon}:=\Gamma^{\varepsilon}(x_{1},y_{1},\dots,x_{n-1},y_{n-1},x_{n})$ (3)
with the additional property that $\varepsilon_{1}\geq\varepsilon_{2}$ implies
$\Gamma^{\varepsilon_{1}}(x_{1},y_{1},\dots,x_{n-1},y_{n-1},x_{n})\subseteq\Gamma^{\varepsilon_{2}}(x_{1},y_{1},\dots,x_{n-1},y_{n-1},x_{n}).$ (4)
This last property is called the property of nested prediction sets.
Recall that $2^{\boldsymbol{Y}}$ is the set of all subsets of $\boldsymbol{Y}$. Thus, a confidence predictor outputs a subset of the label space based on all previous examples and the current object, and the idea is that it should contain the true label $y_{n}$ with a user specified confidence.
We need some more notation. Let $\Gamma$ be a confidence predictor that processes the data sequence
$\omega=(x_{1},y_{1},x_{2},y_{2},\dots)$ (5)
at significance level $\varepsilon$. We say that $\Gamma$ makes an error at the $n$th trial if $y_{n}\notin\Gamma_{n}^{\varepsilon}$. More precisely,
$\text{err}_{n}^{\varepsilon}(\Gamma,\omega):=\begin{cases}1&\text{if $y_{n}\notin\Gamma_{n}^{\varepsilon}$}\\ 0&\text{otherwise,}\end{cases}$ (6)
and the number of errors during the first $n$ trials is
$\text{Err}_{n}^{\varepsilon}(\Gamma,\omega):=\sum_{i=1}^{n}\text{err}_{n}^{\varepsilon}(\Gamma,\omega).$ (7)
2.2 Validity
The number $\text{err}_{n}^{\varepsilon}(\Gamma,\omega)$ is the realisation of a random variable $\text{err}_{n}^{\varepsilon}(\Gamma)$. We say that a confidence predictor $\Gamma$ is exactly valid if, for each $\varepsilon$,
$\text{err}_{1}^{\varepsilon}(\Gamma),\text{err}_{2}^{\varepsilon}(\Gamma),\dots$ (8)
is a sequence of independent Bernoulli random variables with parameter $\varepsilon$. In words, the event of making an error is like getting heads when tossing a biased coin, where the probability of getting heads is always $\varepsilon$. There is also the notion of conservative validity, which is more complicated to state, but essentially means that the error sequence is dominated by a sequence of independent Bernoulli variables with parameter $\varepsilon$. For the complete statement, we refer to [2].
We say that $\Gamma$ is asymptotically valid if
$\lim_{n\to\infty}\frac{1}{n}\text{Err}_{n}^{\varepsilon}(\Gamma)=\varepsilon$ (9)
for each $\varepsilon\in(0,1)$.
2.3 Conformal predictors
Conformal predictors rely on the notion of a nonconformity score. Informally, this is a function that quantifies how “strange” or “unusual” an example $z$ is in relation to what we have seen before. More formally, given a sequence of examples $z_{1},\dots,z_{n}$, we form the bag (or multiset) $\Sigma_{n}:=\lbag z_{1},\dots,z_{n}\rbag$. The set of all bags of size $n$, formed from examples in $\boldsymbol{Z}$ is denoted $\boldsymbol{Z}^{(n)}$. A nonconformity measure is a measurable function
$A:\boldsymbol{Z}^{(*)}\times\boldsymbol{Z}\to\overline{\mathbb{R}}.$ (10)
Definition 2 (Conformal predictor).
The conformal predictor determined by a nonconformity measure $A$ is the confidence predictor defined by setting
$\Gamma^{\varepsilon}(z_{1},\dots,z_{n-1},x_{n}):=\Gamma^{\varepsilon}_{n}:=\left\{y\in\boldsymbol{Y}:\frac{|\{i=1,\dots,n:\alpha_{i}\geq\alpha_{n}\}|}{n}>\varepsilon\right\}$ (11)
where
$\alpha_{i}$ $:=A(\Sigma,(x_{i},y_{i})),$ $i=1,\dots,n-1$ (12)
$\alpha_{n}$ $:=A(\Sigma,(x_{n},y)),$
$\Sigma$ $:=\Lbag(x_{1},y_{1}),\dots,(x_{n-1},y_{n-1}),(x_{n},y)\Rbag.$
The key result about conformal predictors is that they are valid under the exchangeability assumption. A conformal predictor is conservatively valid, and by introducing some randomisation to the special cases when $\alpha_{i}=\alpha_{j}$, which results in a smoothed conformal predictor, we can even get exact validity. However, when exchangeability is violated, these validity guarantees are lost.
The main disadvantage of conformal predictors is their, often, intractable computational cost. In a regression setting, where $\ Y=\mathbb{R}$, we would theoretically have to compute the fraction in (11) for each real number, which is clearly impossible. For some special cases, such as ridge regression, efficient implementations exist, but generally the computational cost is too high. For this reason, inductive conformal predictors (ICP) were introduced. We give a brief description here, and refer the interested reader to [2] for more details. Given a training set of size $l$, we split it into two parts: the proper training set $z_{1},\dots,z_{m}$ of size $m$, and the calibration set $z_{m+1},\dots,z_{l}$ of size $l-m$. For every test object $x_{i}$, compute the prediction set
$\Gamma^{\varepsilon}(z_{1},\dots,z_{l},x_{i}):=\left\{y\in\boldsymbol{Y}:\frac{|\{j=m+1,\dots,l:\alpha_{j}\geq\alpha_{i}\}|+1}{l-m+1}>\varepsilon\right\},$ (13)
where the nonconformity scores are
$\alpha_{j}$ $:=A((z_{1},\dots,z_{m}),z_{j}),$ $j=m+1,\dots,l,$ (14)
$\alpha_{i}$ $:=A((z_{1},\dots,z_{m}),(x_{i},y)).$
These are the most widely used conformal predictors in practice, but as always, there is a price to pay for the computational efficiency. ICPs are no longer exactly valid, but their valid property are of the PAC type, with two parameters. For details on the training-conditional validity of ICPs, see [9].
3 Adaptive conformal inference
Adaptive conformal inference [5] was suggested as a way to achieve asymptotic validity under non-exchangeable data. The idea is that instead of using the same significance level for all predictions, we use the online update
$\varepsilon_{n+1}=\varepsilon_{n}+\gamma(\varepsilon-\text{err}^{\varepsilon_{n}}_{n})$ (15)
where $\gamma$ is a step size (or learning rate). In words, if we made an error in the last step, we increase the significance level, and if not, we decrease it. Gibbs and Candès proved the following finite sample guarantees of ACI.
$\bigg{|}\varepsilon-\frac{1}{N}\sum_{i=1}^{N}\text{err}_{n}^{\varepsilon_{n}}(\Gamma)\bigg{|}\leq\frac{\max\{\varepsilon_{1},1-\varepsilon_{1}\}+\gamma}{\gamma N}~{}(a.s),$ (16)
which in particular converges to 0 as $N\to\infty$, ensuring asymptotic validity. Note that if we want a certain absolute error deviation bound, $\delta>1/n$, for a desired error rate $\varepsilon$, with a finite sample size $N$, we can almost surely achieve it by choosing
$\gamma=\frac{\max\{\varepsilon_{1},1-\varepsilon_{1}\}}{\delta N-1}.$ (17)
The iteration (15) could cause $\varepsilon_{t}\leq 0$ or $\varepsilon\geq 1$. Technically, prediction sets are undefined for confidence predictors, and in particular conformal predictors for $\varepsilon\notin(0,1)$. Thus, we introduce a trivial extension that is compatible with ACI.
Definition 3 (Extended confidence predictor).
An extended confidence predictor is a confidence predictor as defined in definition 1, with the additional property that
$\varepsilon\leq 0$ $\implies\Gamma^{\varepsilon}(x_{1},y_{1},\dots,x_{n-1},y_{n-1},x_{n})=\boldsymbol{Y}$ (18)
$\varepsilon\geq 1$ $\implies\Gamma^{\varepsilon}(x_{1},y_{1},\dots,x_{n-1},y_{n-1},x_{n})=\emptyset.$
Since the difference between confidence predictors and extended confidence predictors is minor, and since any confidence predictor can be trivially modified to be an extended confidence predictor, we will use the terms interchangeably. In analogy with definition 3, we can also define an extended conformal predictor by requiring that the output is $\emptyset$ if $\varepsilon\geq 1$ and $\boldsymbol{Y}$ if $\varepsilon\leq 0$. Similar to confidence predictors and extended confidence predictors, we will use the terms conformal predictor and extended conformal predictor interchangeably.
ACI was introduced in 2021 by Gibbs and Candès [5]. Since then, it has been suggested for time-series forecasting [7], where exchangeability is typically an issue for conformal predictors. It is implemented in this mode in the python package MAPIE [10]. In the time-series domain, it has also been suggested for multi-step ahead forecasting by e.g. [11, 12], and has been shown to be a special case of the conformal PID algorithm [6]. However, despite its name, there is nothing conformal about ACI, as we shall see in the next section.
4 Nothing conformal is required
Interestingly, the proof of (16) does not use any property of a conformal, or even confidence predictor. In fact, all that is required of a predictor $\Gamma$ is that $\emptyset$ is predicted for $\varepsilon\geq 1$ and $\boldsymbol{Y}$ for $\varepsilon\leq 0$. For all other significance levels, even completely random prediction sets ensure (16). The result is proved by Lemma 4.1 in [5], which we reproduce here with notation slightly modified to align with ours.
Lemma 1 (Lemma 4.1 in [5]).
For all $t\in\mathbb{N}$, $\varepsilon_{t}\in[-\delta,1+\delta]$, almost surely, if ACI (15) is used together with a set predictor that outputs $\emptyset$ is predicted for $\varepsilon\geq 1$ and $\boldsymbol{Y}$ for $\varepsilon\leq 0$.
Proof.
Assume that the sequence $\{\varepsilon_{t}\}_{t\in\mathbb{N}}$ is such that $\inf_{t}\varepsilon_{t}<-\gamma$ (the case when $\sup_{t}\varepsilon_{t}>1+\gamma$ is identical). By (15), $\sup_{t}|\varepsilon_{t+1}-\varepsilon_{t}|=\sup_{t}\gamma|\varepsilon-\text{err}_{t}^{\varepsilon_{t}}|<\gamma$. Thus, with positive probability, we may find $t\in\mathbb{N}$ such that $\varepsilon_{t}<0$ and $\varepsilon_{t+1}<\varepsilon_{t}$. However, by assumption, and (15),
$\varepsilon_{t}<0\implies\Gamma_{t}^{\varepsilon_{t}}=\boldsymbol{Y}\implies\text{err}_{t}^{\varepsilon_{t}}=0\implies\varepsilon_{t+1}=\varepsilon_{t}+\gamma(\varepsilon-\text{err}_{t}^{\varepsilon_{t}})\geq\varepsilon_{t}.$ (19)
We have reached a contradiction. ∎
It is clear that no property of conformal, or even confidence predictors, is needed to achieve (16).
5 Why nested prediction sets matter
This section argues that, while not strictly necessary for achieving the finite sample guarantee (16) of ACI, we should restrict ourselves to confidence predictors, as this is the most general type of predictor that satisfy natural assumptions on prediction sets at a confidence level.
5.1 Validity is easy
We have seen that the same guarantee holds for ACI used with a predictor that outputs random subsets of $\boldsymbol{Y}$ for $\varepsilon\in(0,1)$. This predictor is not a confidence predictor because we have dropped the requirement on nested prediction sets. But if we are willing to dispense with nested prediction sets, we could do even better. The coin flip predictor outputs either $\emptyset$ or $\boldsymbol{Y}$ for $\varepsilon\in(0,1)$. Which one is output is determined by flipping a biased coin.
Definition 4 (Coin flip predictor).
A coin flip predictor is a set predictor that for each significance level $\varepsilon\in(0,1)$ outputs the prediction set
$\Gamma_{n}^{\varepsilon}:=\begin{cases}\emptyset,&\text{with probability $\epsilon$}\\ \boldsymbol{Y}&\text{with probability $1-\epsilon$}.\end{cases}$ (20)
The point is that the coin flip predictor is exactly valid. Its error sequence is a realisation of a sequence of independent Bernoulli variables, with parameter $\varepsilon$ for each $\varepsilon\in(0,1)$. However, it is clearly not very informative, and we should be hesitant to use it in practice.
We argue that if we want to predict with confidence, the very least we can ask for is a confidence predictor. Imagine that we play the following game: Before throwing and $n$ sided die, a player is allowed to bet that the outcome will be contained in a subset of ${1,\dots,n}$. Surely, we would not be willing to bet a larger amount on a smaller subset. Dropping the property of nested prediction sets is analogous to this game in the sense that the predictor could state a higher confidence in a smaller prediction set, which is absurd. If we are to mean anything reasonable by stating a confidence level, we simply must require that the prediction sets are nested. For this reason, we argue that, while ACI does not strictly require it, we should restrict its usage to extended confidence predictors in the sense of Definition 3 (and the analogous extended conformal predictors). But that raises a natural empirical question.
5.2 Do we gain anything from conformal predictors?
We have shown that ACI achieves the finite sample guarantee (16) for non-exchangeable data even for non confidence predictors, but we have argued that it would be absurd to drop the property of nested prediction sets. A natural question is whether using ACI with a conformal predictor has any advantages over using it together with a general confidence predictor. For non-exchangeable data, conformal predictors are not necessarily valid, which is also true for confidence predictors.
6 Experiments
We conduct three numerical experiments: one on synthetic data for a regression task, and two on real-world, non-exchangeable datasets, where we tackle both a regression and a classification task. These experiments aim to compare the performance of adaptive conformal inference (ACI) using conformal predictors versus confidence predictors under different data structures.
To evaluate the regression tasks, we employ the mean Winkler interval score (IS) [13], which assesses the quality of the prediction intervals. This score measures both the width of the interval and any penalties for miscoverage. The Winkler interval score for an individual prediction interval is calculated as follows. Let $l$ and $u$ be the lower and upper bounds of the prediction interval for $y$ at the desired significance level $\varepsilon$. The interval score for $[l,n]$ is
$S(l,u,y,\varepsilon):=\begin{cases}(u-l)+\frac{2}{\varepsilon}(l-y),&\text{if $y<l$}\\ u-l,&\text{if $y\in[l,u]$}\\ (u-l)+\frac{2}{\varepsilon}(y-u),&\text{if $y>u$}.\end{cases}$ (21)
The mean Winkler interval score is the mean value of the individual interval scores. Smaller values are better. Since there is a possibility of outputting infinite intervals, we compute interval scores only for finite prediction intervals, and report the fraction of infinite intervals separately. The Winkler interval score is a strictly proper scoring rule [14], meaning in essence that the lowest score is attained if the predicted intervals match the underlying distribution.
In the classification setting, we report the observed excess (OE) of the prediction sets [2], which is defined as
$OE(\Gamma,\varepsilon):=\frac{1}{n}\sum_{i=1}^{n}|\Gamma^{\varepsilon}_{i}\backslash\{y_{i}\},$ (22)
i.e. the average number of false labels included in the prediction sets at significance level $\varepsilon$. It is clear that smaller values are preferable. It is shown in [2] that the observed excess is a conditionally proper efficiency criterion, which in essence mean that the true conditional probability of the label, given the object, is always an optimal conformity score with respect to the conditionally proper efficiency criterion. In this sense, a conditional proper efficiency criterion is analogous to a proper scoring rule. A full discussion on efficiency criteria for conformal classifiers is beyond the scope of this work. For a detailed discussion, see Chapter 3 in [2].
In our evaluation, we have to modify the definition of OE slightly to account for the ACI update (15). We want to achieve the finite sample guarantee (16) with a target error rate $\varepsilon$. Then
$OE_{ACI}(\Gamma,\varepsilon):=\frac{1}{n}\sum_{i=1}^{n}|\Gamma^{\varepsilon_{i}}_{i}\backslash\{y_{i}\}.$ (23)
The difference is that we predict at significance level $\varepsilon_{i}$. This is a natural modification, and we will still refer to it as the observed excess (OE).
6.1 Synthetic data
Our first experiment, based on synthetic data from prior work [4], examines three distinct data settings:
  • 1.
    Setting 1: i.i.d. data. Generate $N=2000$ i.i.d. examples $z_{i}=(x_{i},y_{1})$ with $X_{i}\sim\mathcal{N}(0,I_{4})$ and $Y_{i}\sim X_{i}^{T}\beta+\mathcal{N}(0,1)$ with coefficient vector $\beta=(2,1,0,0)$.
  • 2.
    Setting 2: change points. Same setup as in Setting 1, but the coefficient vector changes. The first 500 points have coefficient vector $\beta_{1}=(2,1,0,0)$. The next 1000 points have coefficient vector $\beta_{2}=(0,-2,-1,0)$, and the last 500 points have coefficient vector $\beta_{3}=(0,0,2,1)$.
  • 3.
    Setting 3: distribution drift. Same setup but the coefficient vector changes from $\beta_{1}=(2,1,0,0)$ to $\beta_{n}(=0,0,2,1)$. The coefficient vector is computed by linear interpolation.
For each setting, we implement the following two methods with the desired error rate set to $\varepsilon=0.1$. For both methods, the significance level for the next prediction is computed using ACI. We ensure that both methods are extended confidence predictors in the sense of Definition 3 by simply outputting $\mathbb{R}$ for $\varepsilon\leq 0$ and $\emptyset$ for $\varepsilon\geq 1$.
  • 1.
    Conformalized least squares (CP). We use the conformalized ridge regression algorithm [2] with ridge parameter set to 0 (which corresponds to least squares). The conformalized ridge regresison algorithm is described in A.1.
  • 2.
    Prediction intervals from ordinary least squares (OLS). The ordinary least squares algorithm can output prediction intervals natively. We implement an online ordinary least squares algorithm, and output the least squares prediction intervals. The resulting confidence predictor is described in A.2.
For all three experiments on synthetic data, the first 100 examples are used as initial training set, and we run them 1000 times using different random seeds (ranging from 0 to 999). In all our experiments, we ask for a target error rate $\varepsilon=0.1$, and set $\varepsilon_{1}=\varepsilon$. Choosing the step size $\gamma$ according to (17) with $\delta=0.05$ would suggest $\gamma\approx 0.0096$.
Figure 1: Numerical simulation for synthetic data showing the average error rate and width of prediction intervals, averaged over 1000 independent trials. Gaps indicate infinite prediction intervals.
IID Changepoints Drift
Method Error Inf IS Error Inf IS Error Inf IS
CP 0.100 0 4.17 0.102 0.00660 9.41 0.104 0.000480 5.63
OLS 0.100 0 4.16 0.102 0.0300 9.28 0.104 0.00155 5.62
Table 1: Numerical simulation for synthetic data showing the average error rate, the fraction of infinite prediction intervals, and the average interval score for the finite intervals, averaged over 1000 independent trials. Inf denotes fraction of infinite prediction intervals, and IS is the mean Winkler interval score.
The results are shown in Figure 1 and Table 1. The first thing to note is that both methods achieve (16) as expected. The gaps in the right-hand side plots in 1 indicate an infinite prediction interval, and the fraction of such infinite intervals is presented in Table 1. We see that although the error rates are identical for both methods, CP outputs fewer infinite intervals in the second and third settings, with change points and drift. As can be seen in Table 1, the difference in terms of average interval score (IS) is not large, but CP outputs fewer infinite intervals on average.
The major difference is the computation time. On average, OLS was about eleven times faster on the same machine. As discussed in A, CP prediction sets can be computed in $\mathcal{O}(n\ln n)$, while $OLS$ require only $\mathcal{O}(n)$.
6.2 Real data
We perform two numerical experiments on publicly available real world datasets. First, a regression task where the aim is to predict the quality of wine, followed by a classification task of handwritten digits.
6.2.1 Wine quality
The Wine Quality dataset is publicly available through the UCI Machine Learning repository [15]. It consists of eleven features that may be useful for predicting the quality of the wine, which is encoded as an integer between 3 and 9, inclusive. Most labels are between 4 and 7. There are 6497 examples in total, 4898 white and 1599 red.
Exchangeability of the wine dataset has been studied in [16], where it is shown that red and white wines are not exchangeable. Thus, we use the white wines as training set, and predict the quality of the red wines.
Predicting the label can be seen either as regression or classification, and we choose to perform the regression task. For confidence predictor, we use the RandomForestQuantileRegressor from the quantile-forest python package [17]. We turn it into an inductive conformal predictor (ICP) by using the WrapRegressor class from the crepes python package [18]. We run our experiment 1000 times using different random seeds for each run. For the ICP, we must set aside a calibration set that is not used in training. Thus, we use 3/4 of the training set as proper training set, and the rest as calibration set. A different random split is used for each run. As a result, the training set for our confidence predictor (QRF) has size 4898 while the ICP has a proper training set size of just 3673 because the remaining data is used for calibration. Again, we choose step size according to (17) with $\delta=0.05$. With $N=1599$ we should choose $\gamma\approx 0.011$.
Figure 2: Numerical experiment on the Wine Quality dataset showing the average error rate and width of prediction intervals, averaged over 1000 independent trials.
Method Error IS Inf
ICP 0.103 5.05 0.00216
QRF 0.0925 4.41 0.0000382
Table 2: Numerical experiment on the Wine Quality dataset showing the mean error rate, interval scores for the finite prediction intervals, and the fraction of infinite intervals output. The results are averaged over 1000 independent trials. Inf denotes fraction of infinite prediction intervals, and IS is the mean Winkler interval score.
Our results are summarised in Figure 2 and Table 2. Again, both methods achieve the theoretical guarantee (16) as expected. Figure 2 summarises the mean running error rate, and the mean interval score (IS), averaged over the 1000 independent trials. It can be noted that CP is overly confident on average early on, while QRF is conservative. In terms of computational efficiency there is not much difference, ICP was introduced to overcome the often intractable computational cost of full CP, and the calibration step is quite fast. The mean error rate together with mean interval scores and the mean fraction of infinite intervals is presented in Table 2. It is worth pointing out that QRF produced infinite intervals at all in only eight out of 1000 trials, while CP did so in 696 trials.
6.2.2 USPS
The US Postal Service (USPS) dataset consists of 7291 training images and 2007 test images of handwritten digits from 0 to 9. The images are $16\times 16$ greyscale pixels. Exchangeability of the USPS dataset has been studied in [19, 20], and it is well known that the examples are not perfectly exchangeable. We use a RandomForestClassifier from the scikit-learn python package [21] with the default parameters, e.g. the forest consists of 100 trees. Again, we use the crepes python package [18], sacrificing 1/4 of the training data for calibration to train an ICP. We also define a confidence predictor by using the predict_proba method, which outputs estimated probabilities for every possible label (details in Appendix B). Choosing step size according to (17) with $\delta=0.05$ and $N=2007$ leads to $\gamma\approx 0.009$.
Figure 3: Numerical experiment on the USPS dataset showing the average error rate and size of prediction sets, averaged over 1000 independent trials.
Method Error Size OE
ICP 0.102 0.944 0.0457
RF 0.0848 0.100 0.0718
Table 3: Numerical experiment on the Wine Quality dataset showing the mean error rate, prediction set size, and observed excess (OE). The results are averaged over 1000 independent trials.
The results of the USPS experiment is summarised in Figure 3 and Table 3. As expected, both methods achieve the theoretical guarantee (16), but RF is quite a bit more conservative than CP, in particular early on, as can be seen in Figure 3. In terms of mean prediction set size, there is not much difference between the methods, but the OE of RF is relatively high, which is likely due to the large prediction sets output by RF in the beginning.
7 Conclusions
In this paper, we have demonstrated that adaptive conformal inference (ACI) does not rely on any specific properties of conformal predictors to achieve its finite sample guarantee (16). In fact, it does not even need the more general concept of a confidence predictor, where the prediction sets are required to be nested (see Definition 1). However, we have argued that the property of nested prediction sets is the very least one should require when predicting with a confidence level. Without it, the coin flip predictor (Definition 4) is exactly valid but rather unhelpful. Since the validity guarantees of conformal predictors are lost if the exchangeability assumption is violated, and ACI provides finite sample coverage guarantees of another kind, we asked if anything is gained by using ACI with a conformal predictor, over using it with a confidence predictor in situations when exchangeability can not be expected.
For our experiments with synthetic data, it seems that there is some advantage to using a conformal predictor insetting 2 and 3 with change points and drift, where the fraction of infinite intervals output by the extended confidence predictor based on ordinary least squares (OLS) was quite a bit larger that for conformalized least squares (CP). The main advantage of OLS lies in its less costly computation. It was about eleven times faster than CP. But these are online algorithms, where all available data is used for training.
When we use an inductive conformal predictor, some part of the training set has to be sacrificed for calibration. In the wine experiment, the confidence predictor RandomForestQuantileRegressor (QRF) performed better than the inductive conformal predictor (ICP) based on the same base algorithm, both in terms of mean interval score, and fraction of infinite intervals. Of 1000 independent trials, only eight contained infinite intervals, while ICP output infinite predictions in 696 trials. In the inductive setting, however, the computational disadvantage of CP disappears, as the calibration step required for ICP is quite fast. It does require sorting, but it only has to be performed once.
For the classification task in the USPS experiment, the picture is less clear. While the ICP based on random forest outperformed the confidence predictor based on the same algorithm in terms of observed excess, it is possible that better results could have been obtained by choosing a larger $\varepsilon_{1}$ to start the ACI iteration (15). It is clear from Figure 3 that the confidence predictor was way too conservative early on, and took some time to adapt. Much of the observed excess, i.e. the number of false labels included in the prediction sets, comes from the first 200 predictions, illustrating that there may be some utility in careful tuning of the initial control input $\varepsilon_{1}$.
We have seen that ACI with a confidence predictor can perform almost as well as with a conformal predictor in the online setting, while being significantly less costly in terms of computation. In the offline mode, a confidence prediction paired with ACI can sometimes outperform an ICP, which may be attributed to that the entire training set can be used to train, without sacrificing some training data for calibration. However, as we have seen, this is not always the case. If the training set is large enough, setting aside some training data for calibration may not be too detrimental. Our experiments are far from conclusive, and more research is needed to determine the relative advantages of ACI with conformal and confidence predictors. Future research should include finding good initial conditions for the control input, and experiments on more complex synthetic data with known properties. Our experiments were limited to using just one underlying machine learning method per dataset. It would be interesting to compare more methods to find out which methods are better suited as confidence predictors, than as ICPs. We speculate that it may come down to how much training data is required to achieve acceptable performance. If less is needed, there is no great loss in sacrificing some data for calibration.
Acknowledgement
The author acknowledge the Swedish Knowledge Foundation and industrial partners for financially supporting the research and education environment on Knowledge Intensive Product Realisation SPARK at Jönköping University, Sweden. Project: PREMACOP grant no. 20220187.
Appendix A Details on the online algorithms
The least squares method goes back to Gauss and Legendre, and is a widely used regression method. Ridge regression is a generalisation of the basic procedure, dating back to the 1960s, introducing a non-negative ridge parameter $a$ (setting $a=0$ recovers ordinary least squares). In matrix form, it can be represented as
$\omega=(X_{n}^{T}X_{n}+aI_{p})^{-1}X_{n}^{T}Y_{n},$ (24)
where $Y_{n}:=(y_{1},\dots,y_{n})^{T}$, $X_{n}=(x_{1},\dots,x_{n})^{T}$, and the ridge regression prediction for an object $x$ is $\hat{y}:=x^{T}\omega$. Lest squares is the special case when $a=0$.
From Eq. (24) we see that the predictions $\hat{y}$ for all objects $x_{i}$ are given by
$\hat{Y}_{n}:=(\hat{y}_{1},\dots,\hat{y}_{n})^{T}=X_{n}(X_{n}^{T}X_{n}+aI_{p})^{-1}X_{n}^{T}Y_{n},$ (25)
or if we define the hat matrix (because it maps $y_{i}$ to its “hatted” form)
$H_{n}=X_{n}(X_{n}^{T}X_{n}+aI_{p})^{-1}X_{n}^{T},$ (26)
we can write $\hat{y}=H_{n}Y_{n}$.
A.1 Conformalised ridge regression
The conformalised ridge regression algorithm (CRR) is a combination of two algorithms; lower CRR and upper CRR, that produce the lower and upper bound respectively. For the upper CRR we use the nonconformity measure $y-\hat{y}$, and the lower uses $\hat{y}-y$, where $\hat{y}$ is the ridge regression prediction of $y$. The complete algorithm is presented in Algorithm 1.
Algorithm 1 Conformalised ridge regression (Alg. 2.4 [2])
Ridge parameter $a\geq 0$, significance level $\varepsilon\in(0,1)$, training set $(x_{i},y_{i})\in\mathbb{R}^{p}\times\mathbb{R},i=1,\dots n-1$ and a test object $x_{n}\in\mathbb{R^{p}}$.
set $C=I_{n}-H_{n}$, $H_{n}$ being defined in Eq. (26)
set $A=(a_{1},\dots,a_{n})^{T}:=C(y_{1},\dots,y_{n-1},0)^{T}$
set $B=(b_{1},\dots,b_{n})^{T}:=C(0,\dots,0,1)^{T}$
for $i=1,\dots,n-1$ do
     if $b_{n}>b_{i}$ then
         set $u_{i}:=l_{i}:=(a_{i}-a_{n})/(b_{n}-b_{i})$
     else
         set $l_{i}=-\infty$ and $u_{i}=\infty$
     end if
end for
sort $u_{1},\dots,u_{n-1}$ in the ascending order obtaining $u_{(1)}\leq\dots\leq u_{(n-1)}$
sort $l_{1},\dots,l_{n-1}$ in the ascending order obtaining $l_{(1)}\leq\dots\leq l_{(n-1)}$
output $[l_{(\lfloor(\varepsilon/2)n\rfloor)},u_{(\lceil(1-\varepsilon/2)n\rceil)}]$ as prediction set.
In our experiments on synthetic data, we trivially extend Algorithm 1 by requiring the output to be (,)($\infty$,$\infty$)( - ∞ , ∞ ) for $\varepsilon\leq 0$ and $\emptyset$ for $\varepsilon\geq 1$, which results in an extended conformal predictor. It is shown in [22] that the prediction sets can be computed in $\mathcal{O}(n\ln n)$ in the online mode. Computing $A$ and $B$ can be done in time $\mathcal{O}(n)$, and sorting can be done in time $\mathcal{O}(n\ln n)$.
A.2 Ridge confidence predictor
The confidence predictor based on ridge regression is summarised in Algorithm 2, where $t_{\varepsilon/2,n-1-p}$ is the critical value from the Student’s $t$-distribution with $n-1-p$ degrees of freedom.
Algorithm 2 Ridge confidence predictor
Ridge parameter $a\geq 0$, significance level $\varepsilon\in(0,1)$, training set $(x_{i},y_{i})\in\mathbb{R}^{p}\times\mathbb{R},i=1,\dots n-1$ and a test object $x_{n}\in\mathbb{R^{p}}$.
set $\hat{y}_{n}=x_{n}^{T}(X_{n-1}^{T}X_{n-1}+aI_{p})^{-1}Y_{n-1}$
set $C=I_{n-1}-H_{n-1}=(c_{1},\dots,c_{n-1})$, $H_{n}$ being defined in Eq. (26)
set $\sigma^{2}=\frac{1}{n-1-p}\sum_{i=1}^{n-1}c_{i}^{2}$
output $[\hat{y}_{n}-t_{\varepsilon/2,n-1-p},[\hat{y}_{n}t_{\varepsilon/2,n-1-p}]$ as prediction set.
Since Algorithm 2 avoids the sorting that is needed in Algorithm 1, it can be computed in $\mathcal{O}(n).$
Appendix B Confidence classifier for the USPS experiment
In the USPS experiment, we turn a random forest classifier from scikit-learn [21] into a confidence predictor. This is achieved by using the predict_proba method, which outputs the predicted label probabilities of an input sample are computed as the mean predicted class probabilities of the trees in the forest. We then have a vector $p=(p_{0},\dots,p_{9})$ of predicted probabilities for the corresponding labels. We include in the prediction set any label $i$ for which $p_{i}>\varepsilon$. Most, but not all, scikit-learn classifiers are equipped with the predict_proba method, and can thus be turned into confidence predictors analogously.
Appendix C More experimental results: Without ACI
For reference, we present the results attained by setting $\gamma=0$, which corresponds to no using ACI at all. The results are summarised in Tables 4-6.
IID Changepoints Drift
Method Error Inf IS Error Inf IS Error Inf IS
CP 0.0982 0 4.15 0.164 0 10.5 0.161 0 5.88
OLS 0.100 0 4.14 0.161 0 10.3 0.162 0 5.86
Table 4: Numerical simulation for synthetic data showing the average error rate, the fraction of infinite prediction intervals, and the average interval score for the finite intervals, averaged over 1000 independent trials.
Method Error Interval score Fraction infinite
ICP 0.210 6.54 0
QRF 0.0661 4.23 0
Table 5: Numerical experiment on the Wine Quality dataset showing the mean error rate, interval scores for the finite prediction intervals, and the fraction of infinite intervals output. The results are averaged over 1000 independent trials.
Method Error Size Observed excess
ICP 0.152 0.863 0.0152
RF 0.0145 1.57 0.581
Table 6: Numerical experiment on the Wine Quality dataset showing the mean error rate, prediction set size, and observed excess (lower values are better for observed excess). The results are averaged over 1000 independent trials.